### v1

# Microscopic and stochastic simulations of chemically active droplets

Abstract
Biomolecular condensates play a central role in the spatial organization of living matter. Their formation is now well understood as a form of liquid-liquid phase separation that occurs very far from equilibrium. For instance, they can be modeled as active droplets, where the combination of molecular interactions and chemical reactions result in microphase separation. However, so far, models of chemically active droplets are spatially continuous and deterministic. Therefore, the relationship between the microscopic parameters of the models and some crucial properties of active droplets (such as their polydispersity, their shape anisotropy, or their typical lifetime) is yet to be established. In this work, we address this question computationally, using Brownian dynamics simulations of chemically active droplets: the building blocks are represented explicitly as particles that interact with attractive or repulsive interactions, depending on whether they are in a droplet-forming state or not. Thanks to this microscopic and stochastic view of the problem, we reveal how driving the system away from equilibrium in a controlled way determines the fluctuations and dynamics of active emulsions.
Introduction.— The formation of biomolecular condensates is a central feature of the spatial organization of living matter at the subcellular and subnuclear levels, and plays a key role in the regulation of multiple metabolic processes Banani et al. (2017); Shin and Brangwynne (2017). During recent years, a significant research effort has been devoted to understanding the physical and chemical mechanisms that govern the formation and the dynamics of these membrane-less organelles, both from experimental and theoretical point of views. The now widely spread picture is that of a form of liquid-liquid phase separation that takes place very far from thermodynamic equilibrium, and that typically results in the selection of a well-defined size for the condensates Hyman et al. (2014); Abbas et al. (2021); Gouveia et al. (2022); Söding et al. (2020).
Figure 1: (a) Two chemical pathways for the interconversions $A\rightleftharpoons B$: a passive pathway, which fulfils detailed balance, and an active pathway, which violates it, because of a constant chemical drive $\Delta\mu$. (b) System under study: chemically active species ($A$ and $B$) and inert particles ($C$, represented as grey dots for readability). See text for details on the interaction potentials. (c) Typical snapshots: starting from the same initial configuration, an equilibrium simulation, where species interconversions only follow the passive pathway ($\Delta\mu=0$), leads to macroscopic phase separation, whereas a nonequilibrium simulation ($\Delta\mu>0$) leads to interrupted phase separation and selection of a finite droplet size.
From a thermodynamical point of view, the stability of finite-size condensates indicates that Oswald ripening is arrested: it was suggested that this could originate from the interplay between phase separation and nonequilibrium chemical reactions  Zwicker (2022). More precisely, in models for active droplets, biomolecular condensates are typically made of proteins that coexist under two states: one in which they tend to form droplets, and one in which they do not. The conversions between these two state are assumed to break detailed balance. This idea, that dates back to the studies of nonequilibrium dissipative structures Glotzer et al. (1994); Christensen et al. (1996), has become prominent in recent theoretical studies and in the context of biomolecular condensates: Flory-Huggins-like continuous descriptions have been designed to account for the formation of active emulsions and chemically active droplets Zwicker et al. (2015, 2017); Weber et al. (2019); Tjhung et al. (2018); Ziethen et al. (2023), and can be engineered experimentally with coarcevates Nakashima et al. (2021, 2018).
However, so far, models of chemically active droplets are spatially continuous and deterministic, in such a way that they cannot account for the polydispersity in size and shape of the droplets. Moreover, these models cannot resolve the dynamics of the system in the stationary state, where fluctuations are responsible for the continuous nucleation of new droplets, and their coalescence. Both these aspects are crucial at the mesoscale, where thermal fluctuations play a prominent role, and which appears to be a particularly relevant level of description for biomolecular condensates Keber et al. (2024). In this context, it is mandatory to design particle-based simulations of chemically active droplets. This represents a challenge, both in the understanding of the relevant timescales and microscopic details of the model, and in the interpretation of the data, which requires advanced computational tools.
In this Letter, we propose a microscopic and stochastic model for chemically active droplets. Using Brownian dynamics simulations of model proteins that switch between two states (Fig. 1(a)) and that transiently attract each other (Fig. 1(b)), we identify the conditions under which stable, finite-size droplets may form. The nonequilibrium ingredient of the model is a chemical drive, which breaks detailed balance, and whose amplitude controls the size of the droplets, their polydispersity, and their shape anisotropy (Fig. 1(c)). In addition, we fully characterize the nonequilibrium steady state, by monitoring the coalescence and shrinkage processes at the level of individual droplets – an aspect out of reach from previous deterministic, coarse-grained description. To our knowledge, the present work constitutes the first particle-based simulations of active droplets that form through the interplay between phase separation and chemical reactions which are driven away from equilibrium in a controlled way.
Model.— We consider a three-dimensional suspension of Brownian particles made of three species, denoted by $A$, $B$ and $C$. $A$ particles may convert into $B$, and vice versa – the dynamics of these interconversions will be specified later on. To account for the high density of the intracellular medium Ellis and Minton (2003); Zimmerman and Minton (1993), $C$ particles are crowders that do not undergo any reaction. Their density is chosen in such a way that the total volume fraction is $0.1$.
We denote by $S_{n}(t)\in\{A,B,C\}$ the species of particle $n$ at time $t$. We assume that the positions of particles $\boldsymbol{r}_{1},\dots,\boldsymbol{r}_{N}$ obey overdamped Langevin equations (see Supplementary Material SM for details on the numerical methods). The pair interaction between two particles $m$ and $n$, denoted by $U_{S_{n}S_{m}}(r_{mn})$, depends on their species and on their relative distance $r_{mn}=|\boldsymbol{r}_{m}-\boldsymbol{r}_{n}|$. The evolution equations of the particles positions read:
$\frac{\mathrm{d}\boldsymbol{r}_{n}}{\mathrm{d}t}=$ $\sqrt{2D}\boldsymbol{\eta}_{n}(t)-\frac{D}{k_{\mathrm{B}}T}\sum_{m\neq n}\nabla U_{S_{n},S_{m}}(r_{mn}),$ (1)
where we assume that all the particles have the same bare diffusion coefficient $D$, and where $\boldsymbol{\eta}_{n}(t)$ is a Gaussian white noise of zero average and unit variance $\langle\eta_{n,i}(t)\eta_{m,j}(t^{\prime})\rangle=\delta_{ij}\delta_{nm}\delta(t-t^{\prime})$. Since the dynamics is overdamped, the velocities of the particles at a given time are irrelevant, and the state of the system is completely described by the configuration vector $\mathcal{C}=(\boldsymbol{r}_{1},\dots,\boldsymbol{r}_{N};S_{1},\dots,S_{N})$. The $B$ particles interact with each other through a Lennard-Jones (LJ) potential, which is truncated at a distance $r_{c}=2.5\sigma$ ($\sigma$ being the diameter of the particles), and shifted in order to ensure continuity of the potential at $r=r_{c}$. It reads $U_{B,B}(r)=[U_{\varepsilon}(r)-U_{\varepsilon}(r_{c})]\theta(r_{c}-r)$, where $U_{\varepsilon}(r)=4\varepsilon\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right]$ is the standard LJ potential and $\theta(r)$ denotes the Heaviside function. All the other pair interactions are purely repulsive and are modeled by the Weeks-Chandler-Andersen (WCA) potential Weeks et al. (1971), which is simply a Lennard-Jones potential truncated and shifted at $r=2^{1/6}\sigma$: $U_{A,\{A,B,C\}}=U_{C,\{A,B,C\}}=[U_{\varepsilon^{\prime}}(r)+\varepsilon^{\prime}]\theta(2^{1/6}\sigma-r)$. The energy parameters of the interaction potentials are $\varepsilon^{\prime}=k_{\mathrm{B}}T$ and $\varepsilon=2k_{\mathrm{B}}T$: the latter ensures that, in the absence of species interconversion, the $B$ particles phase separate at the considered density. Throughout the paper, the distances will be measured in units of $\sigma$, the energies in units of $k_{\mathrm{B}}T$ and times in units of $\sigma^{2}/D$.
In order to specify the rules of species interconversions, let us consider two configurations $\mathcal{C}$ and $\mathcal{C}^{\prime}$, which only differ by the species of one particle. We assume that the species of each particle obey a random telegraph process Gardiner (1985), in such a way that the probability for the system to be in configuration $\mathcal{C}^{\prime}$ at time $t+\delta t$ knowing that it was in configuration $\mathcal{C}$ at time $t$ reads, for a sufficiently small $\delta t$: $P(\mathcal{C}^{\prime},t+\delta t|\mathcal{C},t)\simeq k_{\mathcal{C},\mathcal{C}^{\prime}}\delta t$, where $k_{\mathcal{C},\mathcal{C}^{\prime}}$ is the rate at which the transition takes place. A crucial feature of our model is that the transition rates have two contributions: a passive one $k^{\text{p}}_{\mathcal{C},\mathcal{C}^{\prime}}$, and an active one $k^{\text{a}}_{\mathcal{C},\mathcal{C}^{\prime}}$, in such a way that $k_{\mathcal{C},\mathcal{C}^{\prime}}=k^{\text{p}}_{\mathcal{C},\mathcal{C}^{\prime}}+k^{\text{a}}_{\mathcal{C},\mathcal{C}^{\prime}}$.
First, when the system is at equilibrium, interconversions take place through a ‘passive’ pathway. The conversion rates must obey the detailed balance condition: ${k^{\text{p}}_{\mathcal{C}^{\prime},\mathcal{C}}}/{k^{\text{p}}_{\mathcal{C},\mathcal{C}^{\prime}}}=\exp\{-\beta[E(\mathcal{C})-E(\mathcal{C}^{\prime})]\}$, where $\beta=(k_{\mathrm{B}}T)^{-1}$, and where $E(\mathcal{C})=\frac{1}{2}\sum_{m\neq n}U_{S_{n}S_{m}}(r_{mn})+\sum_{n}w_{S_{n}}$, with $w_{S_{n}}$ being the internal energy of a particle of species $S_{n}$ (we take $w_{A}=w_{C}=0$ and $w_{B}=0.5$). Second, we consider the situation where detailed balance is broken, and where interconversions take place through an ‘active’ pathway. In this situation, we write the ratio between the rates as ${k^{\text{a}}_{\mathcal{C}^{\prime},\mathcal{C}}}/{k^{\text{a}}_{\mathcal{C},\mathcal{C}^{\prime}}}=\exp\{-\beta[E(\mathcal{C})-E(\mathcal{C}^{\prime})+2\kappa_{\mathcal{C}^{\prime},\mathcal{C}}\Delta\mu]\}$ where $\Delta\mu$ is a chemical drive, and $\kappa_{\mathcal{C}^{\prime},\mathcal{C}}=1$ if the transition from $\mathcal{C}^{\prime}$ to $\mathcal{C}$ implied the formation of a $B$ particles and $-1$ otherwise. The active pathway can be interpreted as follows, if we consider for example the common involvement of ATP in biochemical reaction: if the formation of a $B$ particle is only possible by ATP consumption ($A+\text{ATP}\rightleftharpoons B+\text{ADP}$) and assuming that the chemical potentials of ATP and ADP are almost constant (i.e. they are chemostatted), the chemical drive is actually given by the difference between the chemical potentials of the chemostatted species: $\Delta\mu=\mu_{\rm ATP}-\mu_{\rm ADP}$ Jülicher et al. (1997).
To form stable droplets, the formation of $A$ particles must be favored in dense regions (i.e. where droplets tend to form), and, on the contrary, the formation of $B$ particles must be favored in dilute regions. Following the idea from Refs. Zwicker (2022); Zwicker et al. (2017); Weber et al. (2019), we assume that $k^{\text{p}}_{\mathcal{C}^{\prime},\mathcal{C}}=k_{0}(1-\phi_{\text{loc}}/\phi_{\text{max}})\mathrm{e}^{-\frac{\beta}{2}[E(\mathcal{C})-E(\mathcal{C}^{\prime})]}$ and $k^{\text{a}}_{\mathcal{C}^{\prime},\mathcal{C}}=k_{0}(\phi_{\text{loc}}/\phi_{\text{max}})\mathrm{e}^{-\frac{\beta}{2}[E(\mathcal{C})-E(\mathcal{C}^{\prime})+2\kappa_{\mathcal{C}^{\prime},\mathcal{C}}\Delta\mu]}$, where $\phi_{\text{loc}}$ is the local density of $A$ and $B$ particles around the particle whose species change between configuration $\mathcal{C}$ and $\mathcal{C^{\prime}}$, and where $\phi_{\text{max}}$ is the maximum volume fraction of the mixture and is approximated to the maximum packing fraction in 3 dimension: $\phi_{\text{max}}\simeq 0.74$. We fix $k_{0}=10^{-2}$ in all the simulations. With this choice of the reaction rates $k^{\text{a,p}}_{\mathcal{C}^{\prime},\mathcal{C}}$, and at equilibrium ($\Delta\mu=0$), global detailed balance is fulfilled: $k_{\mathcal{C}^{\prime},\mathcal{C}}/k_{\mathcal{C},\mathcal{C}^{\prime}}=\exp\{-\beta[E(\mathcal{C})-E(\mathcal{C}^{\prime})]\}$. Therefore, the parameter $\Delta\mu$ controls finely the deviation from equilibrium, as opposed to previous models of mixtures of particles with ‘active switching’, which have been studied in other contexts Decayeux et al. (2021, 2022, 2023); Alston et al. (2022); Bley et al. (2021); Göth et al. (2022); Bley et al. (2022); Moncho-Jordá and Dzubiella (2020); Longo et al. (2022). In those descriptions, the rates of conversion of the particles are independent of their local environment, in such a way that the distance of the system to equilibrium is difficult to evaluate.
Formation of active droplets.— We perform numerical simulations and tune the parameter $\Delta\mu$, which quantifies the deviation from thermal equilibrium, and which will be varied from 0 to 5. In the range of parameters we consider, we observe that the $B$ particles tend to form droplets (Fig. 1(b,c) and Supplementary Movies). To characterize their spatial structure, we perform a cluster analysis on each trajectory SM , and compute the average volume of the droplets at time $t$, denoted by $\langle v\rangle$.
Figure 2: Left: Average volume of the droplets $\langle v\rangle$ as a function of time for various values of the chemical drive $\Delta\mu$. Solid lines are running averages (over 50 points) of the raw data, shown in lighter color. Right: Radial distribution function (RDF) between $B$ particles in the stationary state (for $t$ between $10^{4}$ and $6\cdot 10^{4}$). The RDF for particles in a system at the same density and interacting via a WCA potential is shown in grey for reference.
At equilibrium, i.e. when $\Delta\mu=0$, we observe that $\langle v\rangle$ increases as a power law ($\langle v\rangle\sim t^{\alpha}$, with $\alpha\simeq 0.48$, see Fig. 2). This slow increase is characteristic of the progressive coarsening of a single droplet, as expected for purely passive systems, through Ostwald ripening. On the contrary, when $\Delta\mu>0$, one observes that $\langle v\rangle$ quickly saturates at a finite value, that depends on $\Delta\mu$, and that is much smaller than the typical volumes reached at equilibrium. We then conclude that the active reaction pathway results in droplet size selection, as predicted by continuous-space reaction-diffusion theories Zwicker et al. (2015, 2017); Zwicker (2022). Note that for values of $\Delta\mu$ smaller than 4, the typical droplet volume that is selected by the active reaction pathway is large, and is likely to be comparable to system size, in such a way that strong finite size effects prevent us from reaching stationary state in a reasonable computational time. For $\Delta\mu>0$, in the stationary state, we compute the radial distribution functions of $B$ particles (Fig. 2). These functions bear the signature of the formation of dense clusters, that become less structured as $\Delta\mu$ increases.
Polydispersity and shape anisotropy of the active droplets.— As seen on Fig. 2, the volume of the droplets strongly fluctuates around its average value. To quantify the polydispersity of the droplets, we measure the volume of each droplet at each time step and in the stationary state, and we plot the resulting histograms on Fig. 3. For small enough values of $\Delta\mu$, the distributions typically show a wide peak at a large value of $v$, which represents the volume of the droplet selected by the active pathway. On this plot, the typical volume reached by a droplet is represented by a star, and denoted $v^{*}$. Furthermore, the distribution of $v$ is large, which shows that the droplets are very polydisperse. This is in contrast with purely deterministic approaches, which result in the selection of a single droplet size, with vanishing fluctuations. Finally, we observe that the variances of these distributions decrease as the chemical drive increases.
Figure 3: Left: Probability distribution of the droplet volume $v$, for various values of the chemical drive $\Delta\mu$, in the stationary state. The two stars represent, for $\Delta\mu=4$ and $4.5$, the maximum of the probability distribution $\mathcal{P}(v)$. Right: Shape anisotropy of the droplets as a function of their volume $v$. Solid lines are running averages (over 10 points), and the widths of the colored areas are the standard deviations (insets: schematics of the spherical and non-spherical droplets, which correspond respectively to low and high values of $\kappa^{2}$).
We also show that polydispersity in volumes is associated with polydispersity in shapes, that we characterize through the computation of the gyration tensors of the droplets and their eigenvalues SM . We introduce a coefficient $\kappa^{2}$, called shape anisotropy, and that varies between 00 and $1$, which correspond respectively to a perfectly spherical droplet and a perfectly elongated one. As seen on Fig. 3, for small enough values of $v$, the shape anisotropy $\kappa^{2}$ decreases with $v$: this is expected, as the $B$ particles constituting the droplets arrange in such a way to minimize surface tension. However, for larger values of $v$, we observe that the shape anisotropy may become an increasing function of the droplets volume. This means that, in this range of parameters, droplets tend to have a more elongated shape than the one that minimizes their surface tension. We interpret this as follows: under the effect of thermal fluctuations, droplets diffuse in the simulation box and coalesce, thus forming large droplets whose shape is elongated (see inset of Fig 3). Under the combined effect of surface tension and active chemical reactions, droplets that are formed by coalescence events relax to smaller and more spherical droplets.
Figure 4: Left: Schematic representation of $\dot{v}$. Right: Phase portrait of the volume of droplets (derivative of the volume with respect to time as a function of the volume) for different values of the chemical drive $\Delta\mu$. For $\Delta\mu=4$ and $4.5$, the two stars correspond that is the typical volume reached by a droplet (shown on Fig. 3).
Size selection mechanism and lifetime of the active droplets.— In order to support this claim, we now study the time evolution of the droplets in the stationary state. Indeed, so far, we have only identified droplets at each timestep without linking them from one step to another. To go further, we design an algorithm to track droplets in time SM . The volume of each droplet is monitored from its birth (that may occur from spontaneous nucleation, or coalescence between two droplets) to its death (that may occur through shrinking below a threshold value, or coalescence with another droplet). For each droplet that is then identified and tracked, we compute the derivative of $v$ with respect to time and plot the resulting phase portrait (the average $\langle\dot{v}\rangle$ as a function of $v$) on Fig 4.
This plot shows that, for values of the chemical drive that lead to the selection of large droplets ($\Delta\mu=4$ or $4.5$), the curves cross the horizontal axis twice: there are then two fixed points. First, at small volumes, the unstable fixed point represented by the empty circle on Fig 4, corresponds to the critical nucleation volume $v_{\rm crit}$: below this volume, droplets are unstable and dissolve spontaneously. Second, the typical volume $v^{*}$ represented by the two stars on Fig. 3 actually corresponds to a stable fixed point. More precisely, when the droplets have initially a volume between $v_{\rm crit}$ and $v^{*}$, the derivative of their volume with respect to time is positive, meaning that they tend to grow by aggregation of $B$ particles from the dilute phase, or coalescence with other droplets. On the contrary, when their initial volume is greater than $v^{*}$, $\langle\dot{v}\rangle$ is negative, meaning that they shrink until they reach the stable value $v^{*}$.
This mechanism for droplet size selection is therefore a microscopic and stochastic counterpart to the predictions from continuous and deterministic models Zwicker et al. (2015). In such theories, the typical volume reached by the droplets at equilibrium is understood as a compromise between a growth (resp. loss) term, which dominates at small (resp. large) volumes. However, as opposed to the ideal case considered in previous theories (droplets in infinite volume without fluctuations), we cannot identify these terms unambiguously given the limited statistics at small volumes.
Finally, we show that our stochastic simulations capture the broad range of behaviors that emerge in the nonequilibrium steady state. To this end, for each droplet identified and tracked in the trajectories, we compute the number of steps elapsed between its birth and its death SM – this ‘lifetime’ will be denoted by $\tau$. We plot on Fig. 5 the probability distributions of $\tau$ for the different values of the chemical drive: they appear to be large, and power-law-like. Interestingly, the average lifetime of the droplets (inset of Fig. 5) depends strongly on the chemical drive, and decreases as the system is taken away from equilibrium. This shows that the active pathway controls not only the size and shape of the droplets, but also the time during which they persist and keep $B$ particles in close vicinity. This property may play a key role in some biological processes, where condensates are known to act as microscopic ‘reactors’, that bypass the diffusion-limited step of chemical reactions Strulson et al. (2012); Castellana et al. (2014).
Figure 5: Distribution of droplet lifetime $\tau$ for fixed $\phi$ at various values of $\Delta\mu$. Average lifetime of the droplets $\langle\tau\rangle$ in the stationary state as a function of $\phi$ for various values of $\Delta\mu$ fig en insert
Conclusion.— In this Letter, we introduced an unprecedented particle-based computational framework designed to study active droplets. Our microscopic and stochastic approach enabled us to simulate Brownian suspensions where Ostwald ripening is interrupted in a controlled way, allowing for a detailed characterization of the polydispersity, shape anisotropy, and fluctuations of active droplets in the nonequilibrium steady state of the system. Our work not only advances the understanding of active emulsions but also provides a versatile methodology applicable to a wide range of systems. As a perspective, the present work opens the way to the understanding of the effect of macromolecular crowding on the formation of biomolecular condensate André and Spruijt (2020); Spruijt (2023).
Acknowledgments.— The authors thank David Zwicker and Chengjie Luo for discussions.